# Quantum noise modeling through Reinforcement Learning

### v2

@@ III.2 Policy
[…]
Figure 2: Schematization of the policy NN for the PPO algorithm.
We have performed different tests to determine the optimal architecture and number of parameters for the NNs used in our policy. These hyperparameters slightly change with the number of qubits. Here, we report the best characteristics observed for circuits with one and three qubits. For the feature extractor, we used a single convolutional layer with 16 filters for single-qubit circuits and 32 filters for three-qubit circuits. This convolutional layer is followed by a dense layer with a ReLU activation function. The optimal number of output features for this dense layer is 64 for single qubit circuits and 32 for three qubits circuits. Both the actor and critic policies are implemented as MLPs with a hidden dense layer containing 256 neurons. The total number of trainable parameters in the entire policy NN is on the order of $10^{4}$. We experimented with increasing the number of parameters by adding additional convolutional layers and increasing the number of features in the feature extractor. However, we observed overfitting when the total number of parameters approached the order of $10^{5}$.
As detailed in Section II.1, in this work we consider a set of four possible noise channels: depolarizing, amplitude damping, and coherent errors $R_{x}$ and $R_{z}$. This means, that the output of our actor policy $\operatorname{A}_{\pi}$ is going to be a $(nqubits,4)$ tensor encoding the predicted parameters of the inserted noise channels at that specific moment. A value of zero corresponds to no noise channel of that type inserted. For example, in a two-qubit circuit, the action
@@ III.3 Training procedure
[…]
Figure 3: Training process of the RL algorithm.
@@
The agent receives the reward at end of each episode and the policy’s NN parameters are then updated to maximize future rewards. The reward is taken to be a function of the distance between the target Density Matrix (DM) ($\rho_{\rm true}$) and the DM of the noisy circuit generated by the agent ($\rho_{\rm agent}$) under some selected metric. We have tested different metrics such as a simple element-wise Mean Squared Error (MSE), density matrix fidelity, and Trace Distance (TD) [62]. Trace distance proved to be the best metric to employ because it is easy to compute and phase invariant. Moreover, it is specifically designed to measure the experimental distinguishability between two quantum states. TD can not be directly used as reward, so we have explored different functional forms. The most effective form has been found to be
$\operatorname{R}(\rho_{\rm agent},\rho_{\rm true})=\frac{1}{\alpha\operatorname{TD}(\rho_{\rm agent},\rho_{\rm true})^{2}+\epsilon}\,,$ (5)
@@ III.4 Dataset Generation
The RL algorithm requires training, testing, and evaluation datasets, which consist of ensembles of random quantum circuits and their corresponding final DMs $\rho_{\rm true}$. These DMs serve as ground truth labels during the training phase of the algorithm (Section III.3). In simulations, the $\rho_{\rm true}$ are computed analytically. However, for circuits executed on hardware, they can be obtained using quantum state tomography [63] or more efficient techniques such as classical shadow state reconstruction [64, 65, 66].
@@
The train set for single qubit circuits is composed of circuits with a fixed number of Clifford gates extracted randomly. Clifford gates are chosen due to, both, their lower simulation cost and their large use in randomized benchmarking and shadow state estimation [52, 68, 65, 66]. In our specific case, since we consider $R_{x}$ and $R_{z}$ gates, the only allowed rotation parameters are multiples of $\pi/2$.
For three-qubit circuits we have used a more sophisticated training set. Half of the training set is composed of circuits with a fixed number of moments with the gates and parameters extracted randomly. The second half of the training set is composed of Clifford circuits implementing randomly chosen three-qubit Clifford unitaries [69]. These circuits do not have a fixed number of gates or moments. We observed an improvement in the generalization properties of the algorithm when using this mixed training set.
@@
For the datasets used in simulations, we defined custom noise models. Specifically, we used two different custom noise models for circuits with one and three qubits. For single-qubit circuits, we applied a depolarizing channel with a depolarizing parameter of $\lambda=0.02$ after each $R_{z}$ gate, and an amplitude damping channel with decay parameter $\gamma=0.03$ after each $R_{x}$ gate. A coherent $R_{x}(\theta^{\prime})$ error with angle $\theta_{x}=0.04\cdot\theta$ is introduced after each $R_{x}(\theta)$ gate. Similarly, a coherent $R_{z}(\theta^{\prime})$ error is added after each $R_{z}(\theta)$ gate, with $\theta_{z}=0.02\cdot\theta$. This noise model is not intended to be realistic but to test our algorithm on a gate dependent noise model. For three-qubit circuits, we used a similar noise on rotation gates and added a depolarizing channel with a depolarizing parameter of $0.02$ and an amplitude damping channel with a decay probability of $0.03$ after each $CZ$ gate. In Section V, we have also used a lower error rate noise model to test the algorithm’s ability to generalize to different noise models. A comprehensive summary of the noise parameters used in the simulations is provided in Table 1.
@@
$CZ$ 0.015 0.01 0 0
Table 1: Noise models used in simulation to train and evaluate the RL algorithm in different conditions. The columns show the noise channels parameters: $\lambda$ for the depolarizing channel, $\gamma$ for the amplitude damping channel and $\theta_{x}$ and $\theta_{z}$ for the coherent error $R_{x}$ and $R_{z}$ respectively. The rows indicate the gates subject to the noise channel, a zero noise parameter means that the noise channel is not applied to that gate.
Generating the training set is straightforward for simulations where the DMs are computed analytically. However, performing full state tomography to obtain these values on quantum hardware can be time-intensive. For this reason, we conducted preliminary tests on one and three qubit simulated circuits to determine the minimal length of the training set. We trained the algorithm on datasets of varying sizes (from $10$ to $10^{3}$), using the noise models introduced in Table 1, and assessed the performance on a common evaluation set. We found that for three qubits, it is sufficient to train the algorithm on a dataset with more than $100$ circuits to avoid overfitting and achieve optimal performance. For single qubit circuits, a dataset with about $20$ circuits is sufficient to obtain nearly optimal performance. This result may vary with the complexity of the noise model; however, it demonstrates that dataset generation is not a bottleneck for quantum chips with a small number of qubits.
@@ IV.1 Simulations
Figure 4: Average density matrix fidelity (left) and trace distance (right) throughout the training process of the RL agent on single-qubit circuits. The metrics have been evaluated on a dataset of 100 circuits using 80% for the train set and 20% for the test set. Error bars report the standard deviation.
Figure 5: Performance evaluation of different noise models on Clifford circuits of depths varying from 3 to 30. The performance is evaluated using average DMs fidelity (left) and trace distance (right). The RL agent has been benchmarked against the RB method, the noiseless scenario and the maximally mixed state (MMS). Error bars report the standard deviation.
@@
Our study begins by training the RL agent to emulate the custom noise model, introduced in Section III.4, on single-qubit circuits. To train the model, we generated a dataset of 100 random Clifford circuits of depth 10. We have used $80\%$ of the dataset for the training, reserving the remaining $20\%$ for the test set. The performance of the agent is evaluated by determining the average fidelity and TD between the DMs it produces and the actual noisy DMs. These values are computed across all episodes for both the training and test sets, as shown in Figure 5. The training process converges after roughly $4\times 10^{5}$ episodes, achieving an average fidelity of about $0.99$ on the test set. The RL agent effectively learns to simulate the noise, exhibiting no signs of overfitting. For the evaluation set, we used a dataset of 100 random non-Clifford circuits with a depth of 15. The average fidelity of the RL agent on the evaluation set is $0.993$, with a standard deviation of $0.003$. This result shows that the agent is able to correctly generalize to non-Clifford circuits.
Figure 6: Average density matrix fidelity (left) and trace distance (right) throughout the training process of the RL agent on three-qubit circuits. The metrics have been evaluated on a dataset of 800 circuits using 80% for the train set and 20% for the test set. Error bars report the standard deviation.
Figure 7: Performance evaluation of different noise models on three-qubit Clifford circuits of depths varying from 3 to 30. The performance is evaluated using average DMs fidelity (left) and trace distance (right). The RL agent has been benchmarked against the RB method, the noiseless scenario and the maximally mixed state (MMS).
@@
To further assess the model’s generalization capability, we evaluate it on random Clifford circuits of varying depths, from 3 to 30. Also in this case we evaluate the performance using the average fidelity and TD between the DMs generated by the model and the actual noisy DMs. Figure 5 compares the performance of the RL agent with the RB noise model described in Section II.1). We also offer a comparison with two limit cases: the DMs of noiseless simulation and the DM of the maximally mixed state (MMS). The RL agent demonstrates its adaptability to circuits of different depths, consistently outperforming RB. This performance suggests that while RB categorizes all noise sources as depolarizing, our algorithm can discern the specific characteristics of the noise. The improvement is especially pronounced on shorter circuits, as the circuit depth increases, the noise approximates to a global depolarizing channel, reducing the relative advantage of our RL agent.
@@
We extended our simulation to three-qubit circuits to evaluate performance in the presence of two-qubit gates. We have used the high noise model on three qubits reported in Table 1. The training set consists of 800 three-qubit circuits, divided into 80% for training (640 circuits) and 20% for testing (160 circuits). This training set is composed of both Clifford and non-Clifford circuits as described in Section III.4. The evolution of the average DMs fidelity and TD throughout the training process is shown in Figure 7. While the algorithm is capable of learning the noise, the convergence is slower than in single-qubit circuits due to a larger action space requiring more episodes for exploration. Convergence is reached after approximately $1.5\times 10^{6}$ episodes achieving an average fidelity of about $0.98$. No signs of overfitting can be observed during the training phase.
For evaluation, we used a dataset of 200 random non-Clifford circuits of depth 15, where the RL agent achieved an average fidelity of $0.98$ with a standard deviation of $0.01$.
Figure 8: Average density matrix fidelity (left) and trace distance (right) throughout the training process of the RL agent on single-qubit circuits executed on a superconductive quantum chip. The metrics have been evaluated on a dataset of 60 circuits using 80% for the train set and 20% for the test set. Error bars report the standard deviation.
@@
Mirroring our previous single-qubit circuit experiments, we evaluated the performance of our RL agent against both the RB method and the limit cases of noiseless circuits and MMS across a variety of circuit depths. This comparison is reported in Figure. 7. In these scenarios, the RL agent consistently demonstrates its adaptability to circuits of different depths, surpassing the performance of the RB method for all circuit lengths.

### v3

@@ III.2 Policy
Observation ($x^{(QCR)}$)CNN features extractor ($\operatorname{FE}$)Features ($x^{(feat)}$)MLP actor ($\operatorname{A}_{\pi}$)MLP critic ($\operatorname{C}_{\pi}$)Rating ($x^{(rate)}$)Action ($x^{(action)}$)
Figure 2: Schematization of the policy neural network for the PPO algorithm. A common feature extractor ($\operatorname{FE}$), composed of a CNN, maps the QCR to a high-dimensional latent feature space. The resulting feature vector serves as input to both the actor policy ($\operatorname{A}\pi$) and the critic policy ($\operatorname{C}\pi$). The actor policy is responsible for selecting the best action for the agent, while the critic policy estimates the future reward for the action chosen by the actor.
We have performed different tests to determine the optimal architecture and number of parameters for the NNs used in our policy. These hyperparameters slightly change with the number of qubits. Here, we report the best characteristics observed for circuits with one and three qubits. For the feature extractor, we used a single convolutional layer with 16 filters for single-qubit circuits and 32 filters for three-qubit circuits. This convolutional layer is followed by a dense layer with a ReLU activation function. The optimal number of output features for this dense layer is 64 for single-qubit circuits and 32 for three-qubit circuits. Both the actor and critic policies are implemented as MLPs with a hidden dense layer containing 256 neurons. The total number of trainable parameters in the entire policy NN is on the order of $10^{4}$. We experimented with increasing the number of parameters by adding convolutional layers and increasing the number of features in the feature extractor. However, we observed overfitting when the total number of parameters approached the order of $10^{5}$.
As detailed in Section II.2, in this work we consider a set of four possible noise channels: depolarizing, amplitude damping, and coherent errors $R_{x}$ and $R_{z}$. This means, that the output of our actor policy $\operatorname{A}_{\pi}$ is going to be a [qubits, 4][\,qubits,\>4\,] tensor encoding the predicted parameters of the inserted noise channels at that specific moment. A value of zero corresponds to no noise channel of that type inserted. For example, in a two-qubit circuit, the action
@@ III.3 Training procedure
Action: place error channelswith chosen parameter.Policy: convolutionalneural network.Environment: quantumcircuit representation.Reward: trace distancebetween density matrices.
Figure 3: Training process of the RL algorithm. At each step the RL agent receives an observation of the environment, the quantum circuit representation of a non-noisy circuit. The agent policy performs an action that consists in putting any number of noise channels at a given circuit moment. At the end of the circuit the reward is computed to minimize the trace distance of the density matrix of the original noisy circuit and the reconstructed one.
@@
The agent receives the reward at end of each episode and the policy’s NN parameters are then updated to maximize future rewards. The reward is taken to be a function of the distance between the target Density Matrix (DM) ($\rho_{\rm true}$) and the DM of the noisy circuit generated by the agent ($\rho_{\rm agent}$) under some selected metric. We have tested different metrics such as a simple element-wise Mean Squared Error (MSE), density matrix fidelity and Trace Distance (TD) [50]. Trace distance proved to be the most suitable metric, as it is both easy to compute and phase-invariant. Moreover, as shown by some of the results presented in Section V, most of the information about the noise is contained in the diagonal elements of the density matrices [42]. Finally, this metric is specifically designed to quantify the experimental distinguishability between two quantum states.TD can not be directly used as reward, so we have explored different functional forms including $1/\operatorname{TD}^{n}$ and $\log(\operatorname{TD})$. The most effective form, as determined through empirical testing, has been found to be
$\operatorname{R}(\rho_{\rm agent},\rho_{\rm true})=\frac{1}{\alpha\operatorname{TD}(\rho_{\rm agent},\rho_{\rm true})^{2}+\epsilon}\,$,
@@ III.4 Dataset Generation
The RL algorithm requires training, testing, and evaluation datasets  [^3] Note that we use the convention more commonly used in the Physics field for naming the datasets, rather than the ML field one. In detail, we call testing set the collection of data used to verify the progress over training, whereas we refer to evaluation set for the data we used to present the final scores obtained by the model., which consist of ensembles of random quantum circuits and their corresponding final DMs ($\rho_{\rm true}$). These DMs serve as ground truth labels during the training phase of the algorithm (Section III.3). In simulations, the $\rho_{\rm true}$ are computed analytically. However, for circuits executed on hardware, they can be obtained using quantum state tomography [17] or more efficient techniques such as classical shadow state reconstruction [85, 26, 15].
@@
The training set for single-qubit circuits is composed of circuits with a fixed number of Clifford gates extracted randomly. Clifford gates are chosen due to, both, their lower simulation cost and their large use in randomized benchmarking and shadow state estimation [38, 48, 26, 15]. In our specific case, since we consider $R_{x}$ and $R_{z}$ gates, the only allowed rotation parameters are multiples of $\pi/2$.For three-qubit circuits we have used a more sophisticated training set. Half of the training set is composed of circuits with a fixed depth with the gates and parameters extracted randomly. The second half of the training set is composed of Clifford circuits implementing randomly chosen three-qubit Clifford unitaries [7]. These circuits do not have a fixed number of gates or fixed depth. We observed an improvement in the generalization properties of the algorithm when using this mixed training set.
@@
For the datasets used in simulations, we defined custom noise models. Specifically, we used two different custom noise models for circuits with one and three qubits. For single-qubit circuits, we applied a depolarizing channel with a depolarizing parameter of $\lambda=0.02$ after each $R_{z}$ gate, and an amplitude damping channel with decay parameter $\gamma=0.03$ after each $R_{x}$ gate. A coherent $R_{x}(\theta_{x})$ error with angle $\theta_{x}=0.04\cdot\theta$ is introduced after each $R_{x}(\theta)$ gate. Similarly, a coherent $R_{z}(\theta_{z})$ error is added after each $R_{z}(\theta)$ gate, with $\theta_{z}=0.02\cdot\theta$. This noise model is not intended to be realistic but to test our algorithm on a gate dependent noise model. For three-qubit circuits, we used a similar noise on rotation gates and added a depolarizing channel with a depolarizing parameter of $0.02$ and an amplitude damping channel with a decay probability of $0.03$ after each $CZ$ gate. In Section V, we have also used a lower error rate noise model to test the algorithm’s ability to generalize to different noise models. A comprehensive summary of the noise parameters used in the simulations is provided in Table 1.
@@
CZCZ 0.015 0.01 0 0
Table 1: Noise models used in simulation to train and evaluate the RL algorithm in different conditions. The columns show the noise channels parameters: $\lambda$ for the depolarizing channel, $\gamma$ for the amplitude damping channel, $\theta_{x}$ and $\theta_{z}$ for the coherent error $R_{x}$ and $R_{z}$ respectively. The rows indicate the gates subject to the noise channel, a zero noise parameter means that the noise channel is not applied to that gate.
Generating the training set is straightforward for simulations where the DMs are computed analytically. However, performing full state tomography to obtain these values on quantum hardware can be time-intensive. For this reason, we conducted preliminary tests on one- and three-qubit simulated circuits to determine the minimal length of the training set. We trained the algorithm on datasets of varying sizes (from $10$ to $10^{3}$), using the noise models introduced in Table 1, and assessed the performance on a common evaluation set. We found that for three qubits, it is sufficient to train the algorithm on a dataset with more than $100$ circuits to avoid overfitting and achieve optimal performance. For single-qubit circuits, a dataset with about $20$ circuits is sufficient to obtain nearly optimal performance. This result may vary with the complexity of the noise model; however, it demonstrates that dataset generation is not a bottleneck for quantum chips with a few qubits.
@@ IV.1 Simulations
Figure 4: Average density matrix fidelity throughout the training process of the RL agent on single-qubit circuits (left) and three-qubit circuits (right). The metrics were evaluated on a dataset of 100 circuits for the single-qubit case and 800 circuits for the three-qubit case, using 80% for the training set and 20% for the test set. Shaded regions represent the standard deviation.In the single-qubit case, convergence of the training process was reached after approximately $4\times 10^{5}$ episodes, achieving an average fidelity of about $0.99$ on the test set. For the three-qubit case, convergence was reached after approximately $1.5\times 10^{6}$ episodes, achieving an average fidelity of about $0.98$ on the test set. No significant overfitting was observed in either case.
Figure 5: Performance evaluation of different noise models on Clifford circuits with depths varying from 3 to 30 for the single-qubit case (left) and the three-qubit case (right). Performance is measured using average density matrix fidelity; shaded regions represent the standard deviation. The RL agent was benchmarked against the RB method, the noiseless scenario, and the maximally mixed state (MMS). The absolute advantage of using the RL model is evident when compared to the case with no noise channel insertion. Moreover, in both cases, the RL model outperforms the RB model. Fidelity between the noisy circuit and the MMS is provided as a reference for the overall noise level; as the depth increases, the circuit’s behavior can approximate that of a completely depolarizing channel, converging toward the MMS.
@@
Our study begins by training the RL agent to emulate the custom noise model, introduced in Section III.4, on single-qubit circuits. To train the model, we generated a dataset of 100 random Clifford circuits of depth 10. We have used $80\%$ of the dataset for the training, reserving the remaining $20\%$ for the test set. The performance of the agent is evaluated by determining the average fidelity between the DMs it produces and the actual noisy DMs. These values are computed across all episodes for both the training and test sets, as shown in Figure 5. The training process converges after roughly $4\times 10^{5}$ episodes, achieving an average fidelity of about $0.99$ on the test set. The RL agent effectively learns to simulate the noise, exhibiting no signs of overfitting. For the evaluation set, we used a dataset of 100 random non-Clifford circuits with a depth of 15. The average fidelity of the RL agent on the evaluation set is $0.993$, with a standard deviation of $0.003$. This result shows that the agent is able to correctly generalize to non-Clifford circuits. In this kind of circuits, every possible rotation angle is allowed, while the training set consists of Clifford circuits with fixed rotation angles multiples of $\pi/2$. A summary of the performance obtained by the model during the training and testing is reported in Table 2.
@@
To further assess the model’s generalization capability, we evaluate it on random Clifford circuits of varying depths, from 3 to 30. Also in this case we evaluate the performance using the average fidelity between the DMs generated by the model and the actual noisy DMs. Figure 5 compares the performance of the RL agent with the RB noise model described in Section II.2). We also offer a comparison with two limit cases: the DMs of noiseless simulation and the DM of the maximally mixed state (MMS). The noiseless simulation serves as a baseline, representing an agent that does not apply any noise channels, and thus provides a reference for evaluating the benefit of using the RL agent. The maximally mixed state (MMS) is included as an additional reference to indicate the total noise level; as circuit depth increases, the system’s behavior can approximate that of a completely depolarizing channel, converging toward the MMS. The RL agent demonstrates its adaptability to circuits of different depths, consistently outperforming RB. This performance suggests that while RB categorizes all noise sources as depolarizing, our algorithm can discern the specific characteristics of the noise.
Training set Test set Evaluation set Clifford depth 30
1 qubit $9\pm 0.01 0.99±0$$9\pm 0.01$ $93\pm 0.003 0.99±0$$9\pm 0.01$
3 qubits $0.98\pm 0.02$ $0.98\pm 0.02$ $0.98\pm 0.01$ $0.98\pm 0.01$
Table 2: . Summary of the RL model performance across different training, test, and evaluation sets for single- and three-qubit circuits. Performance is reported as the average fidelity between the reconstructed and original density matrices, with standard deviation as the error. The training and test sets consist of Clifford circuits for the single-qubit case, and both Clifford and non-Clifford circuits for the three-qubit case. The evaluation set includes non-Clifford circuits of depth 15 in both cases. Fidelity on Clifford circuits of depth 30 is also reported as an alternative evaluation set.
@@
We extended our simulation to three-qubit circuits to evaluate performance in the presence of two-qubit gates. We have used the high noise model on three qubits reported in Table 1. The training set consists of 800 three-qubit circuits, divided into 80% for training (640 circuits) and 20% for testing (160 circuits). This training set is composed of both Clifford and non-Clifford circuits as described in Section III.4. The evolution of the average DMs fidelity throughout the training process is shown in Figure 5. While the algorithm is capable of learning the noise, the convergence is slower than in single-qubit circuits due to a larger action space requiring more episodes for exploration. Convergence is reached after approximately $1.5\times 10^{6}$ episodes achieving an average fidelity of about $0.98$. No signs of overfitting can be observed during the training phase.For evaluation, we used a dataset of 200 random non-Clifford circuits of depth 15, where the RL agent achieved an average fidelity of $0.98$ with a standard deviation of $0.01$.
@@
Mirroring our previous single-qubit circuit experiments, we evaluated the performance of our RL agent against both the RB method and the limit cases of noiseless circuits and MMS across a variety of circuit depths. This comparison is reported in Figure 5. In these scenarios, the RL agent consistently demonstrates its adaptability to circuits of different depths, surpassing the performance of the RB method for all circuit lengths.A summary of the RL model performance on training, test, and evaluation sets is reported in Table 2.
Figure 6: Distribution of the noise channels parameters, depolarizing parameter (lambda) top and the amplitude damping parameter (gamma) bottom, inserted by the RL agent on a dataset of 100 random three-qubit Clifford circuits of depth 15. The average values of lambda and gamma are respectively $0.018$ and $0.029$, which are very close to the real parameters of the noise model used in the simulation ($0.02$ and $0.03$ respectively).
In the results presented in this section we have shown that the RL agent is able to reproduce the noise model used in the simulation with high accuracy. The limitation of this algorithm, which is a common limitation of ML algorithms, is that it is difficult to extrapolate information about the noise model. Learning information about specific features of the noise may be useful to improve the reliability of the chip, for example by addressing the most problematic gates. However, the RL agent behaves like a “black box” making it difficult to understand the reasons behind its decisions. A deep understanding of the agent behavior would require the application of explainable ML techniques [1, 67, 39], which is beyond the scope of this work. In order to have a few insights about the noise model learned by the RL agent, we can keep track of the number of noise channels and parameter values inserted by the agent. For this reason we have generated a dataset composed of 100 random three-qubit Clifford circuits of depth 15, and we have applied the high noise model reported in Table 1 to these circuits. We have applied the RL algorithm to this dataset, and we have recorded the number of depolarizing and amplitude damping channels inserted by the agent and their parameters. Figure 6 shows the distribution of the parameters of these noise channels. It is interesting to observe that, while the distributions are not peaked around the real parameters of the noise model, the average values of the depolarizing parameter (lambda) and the amplitude damping parameter (gamma) are respectively $0.018$ and $0.029$, which are very close to the real parameters of the noise model used in the simulation ($0.02$ and $0.03$ respectively). The distributions of the parameters lambda and gamma are significantly different. The former is peaked around lower values and the shape is similar to a Poisson distribution, while the latter has a more uniform shape. Moreover, we can observe that the distribution of the amplitude damping parameter has a peak on the maximum value of $0.06$. This means that the agent is constrained by the maximum value of the parameter. The average number per circuit of noise channels inserted by the agent is $19.7$ for the depolarizing channel and $20.2$ for the amplitude damping channel. Also, this number is coherent with the noise model used in the simulation, where, for random generated circuits of depth 15, we expect to have about $20$ depolarizing and amplitude damping channels per circuit. This study shows that we can extract some average noise parameters from the RL agent by looking at the number of inserted noise channels and their parameters. However, this does not provide a complete understanding of the noise model learned by the agent and learning specific features of the noise. The application of explainable ML techniques to this algorithm is a possible future direction of this work.